At noon a private plane left Austin for Los Angeles, 2100 km away, flying at 500 km/hr. One hour later, a jet left Los Angeles for Austin at 700 km/hr. What time did they pass each other?
Answer by jorel1380(2518)

500+500x+700x=2100
1200x=1600
x=1600/1200=4/3 hours=1 hr. 20 min.
1 p.m. + 1 hr. 20 min.=2:20 p.m.